# Imbalance in Balance: Online Concept Balancing in Generation Models

### v1

@@ Abstract
In visual generation tasks, the responses and combinations of complex concepts often lack stability and are error-prone, which remains an under-explored area. In this paper, we attempt to explore the causal factors for poor concept responses through elaborately designed experiments. We also design a concept-wise equalization loss function (IMBA loss) to address this issue. Our proposed method is online, eliminating the need for offline dataset processing, and requires minimal code changes. In our newly proposed complex concept benchmark Inert-CompBench and two other public test sets, our method significantly enhances the concept response capability of baseline models and yields highly competitive results with only a few codes.
@@
00footnotetext: Corresponding author.
@@ 1 Introduction
In recent years, visual generative models have witnessed remarkable progress, attributed to the improvements of generation paradigms (e.g. diffusion models  [47, 28], auto-regressive models [25, 49, 23, 48, 56]), the proposal of new network architectures (e.g. DiT [37, 12, 6], etc.), and the release of large-scale dataset (e.g. [45, 8, 51], etc.). Driven by the concerted efforts of academia, industry, and the community, these models have reached a high level of sophistication. They are now capable of generating strikingly realistic images ([44, 43, 38, 2, 33, 12]), videos ([22, 17, 32, 3, 30]), and 3D models ([26, 57, 54]), and have been extensively adopted in creative endeavors and content generation scenarios. Remarkably, the most attractive aspect of generative models is their proficiency in deciphering and recombining real-world concepts. This enables them to conjure up objects and scenes that do not even exist in the physical world. This distinctive capability significantly expands the creative boundaries for content producers, unlocking novel opportunities for innovation.
However, in real-world applications, generative models often struggle to consistently generate outputs that closely match user expectations, especially in the realm of concept composition. Consider text-to-image (T2I) generation: even the most advanced models [2, 12, 33, 14] often suffer from problems like concept missing, attribute leakage and concept coupling as shown in Figure 1. To address these issues, several training-free approaches [5, 29, 52, 7, 13] have been proposed. Although these methods can be effective in specific scenarios, we contend that several essential problems underlying this challenge still remain under-explored:
@@
Visual generation models are complex systems with multiple components, and their performance is affected by many factors. Previous efforts [34, 53, 60] have yielded interesting insights, often using simple synthetic data (e.g. basic shapes of different sizes and colors) or focusing on class-to-image (C2I) tasks. However, these data are too simplistic to mirror the true complexity of text-to-image (T2I) tasks, where the number of concepts and their combinations far exceed those studied before. To bridge this gap, we are the first to conduct in-depth analysis on large-scale text-image pair data. We propose and verify the following hypotheses one by one: 1) When the dataset reaches a sufficient scale, concepts will naturally be covered and learned comprehensively. 2) As the model size increases, it becomes easier for the model to learn concept responses effectively. 3) The distribution of the dataset itself plays a dominant role. Through elaborately designed experiments, we have uncovered some intriguing findings. Firstly, an increase in dataset scale does not lead to improved responses for combined concepts. Secondly, once the model size reaches a certain threshold, there is no further enhancement in complex concept responses. Thirdly, a more balanced data distribution can significantly boost the model’s ability to respond to combined concepts.
@@
Based on previous analysis, our work, like prior studies [24, 19], focuses on optimizing data distribution. However, this endeavor is rife with challenges. First, the large data scale in text-to-image (T2I) tasks renders any preprocessing for dataset-level statistical analysis and equalization [55, 21] prohibitively costly. Second, since each text prompt typically contains multiple concepts, sample equalization using loss weights [9, 36] has limited effectiveness in balancing concept distributions. Third, given the diverse generation paradigms and high model training costs, an ideal solution should be plug-and-play and applicable across different models.
@@
Unlike general image generation benchmarks, the ability to combine concepts can only be differentiated on more complex and targeted test sets. In addition to evaluating on existing benchmarks such as T2I-CompBench [18] and LC-Mis benchmark [60], we have carefully constructed a new benchmark called Inert-CompBench. We identify inert concepts (difficult to integrate with other concepts) from large-scale text-image datasets and combine them with head concepts to obtain caption candidates. Then we construct a concept graph and filter out the uncommon-sense compliant concept pairs based edge weight. Finally, we generate 5 captions for each pairs with LLM [1] for evaluation.
@@ 2.1 Concept Composition
Concept composition is the ability of generative models to accurately generate content with multiple concepts, reflecting their learning and understanding capabilities. It is a crucial indicator of a model’s generalization ability. In text-image generation, several benchmarks [5, 18, 60] have been proposed to comprehensively evaluate this ability. However, as shown in Figure 1, existing pre-trained models [12, 33, 2, 38] still suffer from issues like missing object, attribute leakage and concept entanglement. Some studies [34, 4, 53] based on synthetic experiments find that the concept composition ability of diffusion models is related to data completeness, balance, and disentanglement. However, these studies are often small-scale and class-conditioned, creating a gap with text-image tasks. From an application perspective, some works [5, 13, 29, 52] propose training-free methods by optimizing attention maps to enhance the model’s concept response strength, when others [31, 11] add input modalities or generating through multiple rounds of feedback. These methods are often limited by the capabilities of the foundation model. And there is little work analyzing the factors determining concept composition ability from a pre-training perspective. Our work aims to address these gaps.
@@ 2.2 Data Balancing
Data balancing aims at learning generalized models from long-tailed data distributions. Many works [59, 24, 19] have achieved excellent results in class-specific tasks, such as re-sampling, re-margining, and re-weighting [27, 9, 36, 42]. Some studies [39, 55, 50, 58, 21] have extended class-based re-weighting to class-image generation tasks with inter-class distance, balanced distribution, label augmentation, or self-guided methods, showing impressive results. However, since text prompts is a joint distribution of multiple classes and each image cannot be assigned a single class, data balancing in text-image tasks still needs further exploration.
@@ 3 Causal Factors of Concept Composition
In this section, we bridge the gap between synthetic experiments [34, 53, 4] and text-image generation tasks to further explore the casual factors of concept composition. We conduct controlled experiments on text-image datasets and investigate the influence of crucial factors: model size, dataset scale, and data distribution.
@@
Experiment Setup To ensure dataset the same and training efficiency, we re-caption open-source datasets [45] and collect a higher quality dataset, as existing open-source pre-trained models [33, 12, 2] do not provide their training data. Our dataset contains 31M text-image pairs with an average text length of 100 words and 20K noun concepts with distributions shown in Figure 2. To ensure the generality of the conclusion, we employ a DiT-based diffusion model [37] without any special design. To accurately evaluate the model’s concept composition ability rather than clarity, we introduce VQA based on VLM [35] for success rate as a quantitative metric except the CLIP-Score [40]. We detect all noun concepts in the caption and pair them, requiring the VLM to verify the existence and relationships of concepts with the following questions: ”Is concept A present in the image?”, ”Is concept B present in the image?” and ”Does the relationship between concepts A and B match the caption?”. When all responses are ”yes,” the sample is considered successful.
Figure 3: The performance of models with different parameter sizes under the LC-Mis benchmark [60].
Model Size In generation tasks, model size also follows the scaling law [20], where models with more parameters have stronger generation ability. We keep the VAE from Stable Diffusion [12] the same and train diffusion models with 100M, 200M, 500M and 1B parameters on the same 31M dataset from scratch, differing only in the number of blocks and channels. We evaluate the CLIP score and success rate on the LC-Mis benchmark, as shown in Figure 3. When the model size exceeds 200M, the concept composition ability increases at a much slower rate as the model size grows.It indicates that model size is no longer the casual factor for concept composition when it reaches a certain magnitude relative to the dataset.
@@ 4.1 Theoretical Analysis
Without loss of generality, we derive with the $\epsilon$-prediction DDPM [16] framework. It is worth noting that this approach can be easily extended to various other diffusion model variants, such as flow matching [28]. In common implementations [10, 47], the training loss of diffusion models parameterized by $\theta$ can be written as follow:
@@
where $y$ is the text prompts, $x_{t}$ is the noisy image at timestep $t$ from image $x_{0}$ and random Gaussian noises $\epsilon$. Inspired by existing work [46] where image regions respond to prompts as concept phrases, we rewrite the loss function as follow:
@@
On the other hand, we wish that the model responds with similar intensity to different concepts [5]. Thus, we have the response intensity represented by the difference between the conditional and unconditional distribution as follow:
@@
where the intensity depends only on $a_{t}$. Meanwhile, we have the formulation of conditional and unconditional distribution in classifier-free guidance [15] as follows:
@@
$=\left\|\frac{\sigma_{t}}{\sum_{i=1}^{n}\varphi(c_{i})}\left[\nabla_{a_{t}}\log p\left(c_{j}|a_{t}\right)\sum_{i=1,i\neq j}^{n}\varphi(c_{i})\right.\right.$
@@ 4.2 IMBA Distance
Measuring data balance is particularly challenging due to the complexity of text prompts compared to classes [27, 39]. Additionally, with the exponential growth of datasets, offline data pruning will lead to data waste and significant computational and time costs. In this section, we demonstrate that the IMBA distance is able to represent the frequency proportion of concepts with both synthetic experiments and text-image generation experiments. We further give out the formulation of IMBA distance in Appendix A.
@@ 4.3 IMBA Loss
Compared to offline concept frequency-based loss weights [59, 9, 36], IMBA loss is more accurate and efficient. Firstly, since text prompts are the joint distribution of multiple concepts, and each concept has a different frequency, it is challenging to intuitively and quickly set loss weights for samples as in class-related tasks. Secondly, while it is feasible to construct a concept graph where concepts are represented as nodes and the co-occurrence frequency of concept pairs as edge values to derive loss weights, the data distribution that the model learns evolves throughout the training process. This evolution creates a misalignment between the offline loss weights and the model’s understanding. In contrast, IMBA distance is naturally coupled with the training process, providing a more consistent representation of data distribution. Thirdly, since different regions of an image may contain different concepts, data balancing should be performed at the concept region level. IMBA loss supports different loss weights for different regions, making it more accurate compared to a single weight for all regions. Additionally, with the exponential growth of data, constructing an offline concept graph becomes increasingly time-consuming and computationally intensive, making it challenging to reuse across different datasets. IMBA loss, on the other hand, only requires the calculation of IMBA distance with only a few lines of code during training, making it highly efficient and easily reusable. It is worth noting that our method is orthogonal to training-free methods and can be combined to further enhance the concept composition ability.
@@ 4.4 Inert-CompBench
Due to the lack of attention to these concepts in existing benchmarks [18, 60], we extract inert concepts from open-world datasets to construct a new benchmark called Inert-CompBench as a supplement. As shown in Algorithm 2, our framework comprises five phases: (1) Extract candidate sets where head concepts exhibit occurrence frequencies exceeding tail concepts by 100:1 ratio based on large-scale dataset statistics. (2) Select n domain-representative entity concepts from each pool through semantic typicality analysis. (3) Construct n×n Cartesian product combination space. (4) Build concept co-occurrence graphs to filter Top-K pairs with minimal structural associations, ensuring test cases reflect non-trivial compositional relationships. (5) Generate 5 linguistically diverse prompts per selected pair using GPT-4, ultimately forming a benchmark with 1K fine-grained test instances. This design forces models to handle statistically weak concept combinations, effectively revealing their compositional reasoning limitations.
@@
3:Select $n$ representative entities: ^$\hat{\mathcal{H}}\subset\mathcal{H},\hat{\mathcal{T}}\subset\mathcal{T}$ via semantic typicality analysis (Phase 2)
@@
7:Benchmark set =$\mathcal{B}=\{\mathcal{P},\mathcal{C}^{*}\}$
@@ 5 Experiments
Setup. To keep the dataset the same and improve training efficiency, we recaption and filter open-source data [45], resulting in a higher quality dataset of 31M samples for training. We employ a 1B parameter DiT-based diffusion model [37] without any special design as the pipeline. Text prompts are injected into the diffusion model using T5 model [41], and $512\times 512$ images are encoded into the latent space using the VAE from Stable Diffusion [12].
Baseline. We train the diffusion model from scratch for 4 epochs on the new dataset using the original diffusion loss to establish a baseline. To demonstrate the effectiveness of our method, we train the same model for the same epochs on the same dataset with the proposed IMBA loss. Additionally, we implement the training-free method [5] on the baseline for comparison. Furthermore, we finetune the baseline model from 3 epochs for 1 additional epochs with the IMBA loss, demonstrating that our method also provides significant benefits during fine-tuning.
Benchmark. We evaluate the concept composition ability of models on T2I-CompBench [18], LC-Mis [60] and Inert-CompBench. We employ VQA based on VLM [35] for evaluation in LC-Mis and Inert-CompBench as described in experiments setup of Section 3.
@@
Model LC-Mis [60] T2I-CompBench [18] Inert-CompBench

### v2

@@ Abstract
In visual generation tasks, the responses and combinations of complex concepts often lack stability and are error-prone, which remains an under-explored area. In this paper, we attempt to explore the causal factors for poor concept responses through elaborately designed experiments. We also design a concept-wise equalization loss function (IMBA loss) to address this issue. Our proposed method is online, eliminating the need for offline dataset processing, and requires minimal code changes. In our newly proposed complex concept benchmark Inert-CompBench and two other public test sets, our method significantly enhances the concept response capability of baseline models and yields highly competitive results with only a few codes released at https://github.com/KwaiVGI/IMBA-Loss.
@@
00footnotetext: *Done during internship at Kling Team. Corresponding author.
@@ 1 Introduction
In recent years, visual generative models have witnessed remarkable progress, attributed to the improvements of generation paradigms (e.g. diffusion models  [song2020score, lipman2022flow], auto-regressive models [mar, var, lee2022autoregressive, sun2024autoregressivebeats, titok]), the proposal of new network architectures (e.g. DiT [dit, sd3, chen2023pixart], etc.), and the release of large-scale dataset (e.g[laion, chen2024panda, wang2024koala], etc.). Driven by the concerted efforts of academia, industry, and the community, these models have reached a high level of sophistication. They are now capable of generating strikingly realistic images ([imagen, latentdiffusion, podell2023sdxl, dalle3, midjourney, sd3]), videos ([kling, imagenvideo, snapvideo, sora, latte]), and 3D models ([li2024craftsman, zhang2024clay, wu2024direct3d]), and have been extensively adopted in creative endeavors and content generation scenarios. Remarkably, the most attractive aspect of generative models is their proficiency in deciphering and recombining real-world concepts. This enables them to conjure up objects and scenes that do not even exist in the physical world. This distinctive capability significantly expands the creative boundaries for content producers, unlocking novel opportunities for innovation.
However, in real-world applications, generative models often struggle to consistently generate outputs that closely match user expectations, especially in the realm of concept composition. Consider text-to-image (T2I) generation: even the most advanced models [dalle3, sd3, midjourney, flux] often suffer from problems like concept missing, attribute leakage and concept coupling as shown in Figure 1. To address these issues, several training-free approaches [chefer2023attend, liu2022compositional, wang2023compositional, chen2024traininglayout, feng2022trainingstructure] have been proposed. Although these methods can be effective in specific scenarios, we contend that several essential problems underlying this challenge still remain under-explored:
@@
Visual generation models are complex systems with multiple components, and their performance is affected by many factors. Previous efforts [okawa2023compositionalemerge, wiedemer2023compositionalprinceples, zhao2024lost] have yielded interesting insights, often using simple synthetic data (e.g. basic shapes of different sizes and colors) or focusing on class-to-image (C2I) tasks. However, these data are too simplistic to mirror the true complexity of text-to-image (T2I) tasks, where the number of concepts and their combinations far exceed those studied before. To bridge this gap, we are the first to conduct in-depth analysis on large-scale text-image pair data. We propose and verify the following hypotheses one by one: 1) When the dataset reaches a sufficient scale, concepts will naturally be covered and learned comprehensively. 2) As the model size increases, it becomes easier for the model to learn concept responses effectively. 3) The distribution of the dataset itself plays a dominant role. Through elaborately designed experiments, we have uncovered some intriguing findings. Firstly, an increase in dataset scale does not lead to improved responses for combined concepts. Secondly, once the model size reaches a certain threshold, there is no further enhancement in complex concept responses. Thirdly, a more balanced data distribution can significantly boost the model’s ability to respond to combined concepts.
@@
Based on previous analysis, our work, like prior studies [leevy2018surveyimbalance, johnson2019surveyclassimbalance], focuses on optimizing data distribution. However, this endeavor is rife with challenges. First, the large data scale in text-to-image (T2I) tasks renders any preprocessing for dataset-level statistical analysis and equalization [yan2024trainingclassimbalance, katsumata2024labelaug] prohibitively costly. Second, since each text prompt typically contains multiple concepts, sample equalization using loss weights [cui2019classloss, park2021influence] has limited effectiveness in balancing concept distributions. Third, given the diverse generation paradigms and high model training costs, an ideal solution should be plug-and-play and applicable across different models.
@@
Unlike general image generation benchmarks, the ability to combine concepts can only be differentiated on more complex and targeted test sets. In addition to evaluating on existing benchmarks such as T2I-CompBench [huang2023t2icompbench] and LC-Mis benchmark [zhao2024lost], we have carefully constructed a new benchmark called Inert-CompBench. We identify inert concepts (difficult to integrate with other concepts) from large-scale text-image datasets and combine them with head concepts to obtain caption candidates. Then we construct a concept graph and filter out the uncommon-sense compliant concept pairs based edge weight. Finally, we generate 5 captions for each pairs with LLM [achiam2023gpt] for evaluation.
@@ 2.1 Concept Composition
Concept composition is the ability of generative models to accurately generate content with multiple concepts, reflecting their learning and understanding capabilities. It is a crucial indicator of a model’s generalization ability. In text-image generation, several benchmarks [chefer2023attend, huang2023t2icompbench, zhao2024lost] have been proposed to comprehensively evaluate this ability. However, as shown in Figure 1, existing pre-trained models [sd3, midjourney, dalle3, podell2023sdxl] still suffer from issues like missing object, attribute leakage and concept entanglement. Some studies [okawa2023compositionalemerge, chang2024skews, wiedemer2023compositionalprinceples] based on synthetic experiments find that the concept composition ability of diffusion models is related to data completeness, balance, and disentanglement. However, these studies are often small-scale and class-conditioned, creating a gap with text-image tasks. From an application perspective, some works [chefer2023attend, feng2022trainingstructure, liu2022compositional, wang2023compositional] propose training-free methods by optimizing attention maps to enhance the model’s concept response strength, when others [manas2024improving, ding2024freecustom] add input modalities or generating through multiple rounds of feedback. These methods are often limited by the capabilities of the foundation model. And there is little work analyzing the factors determining concept composition ability from a pre-training perspective. Our work aims to address these gaps.
@@ 2.2 Data Balancing
Data balancing aims at learning generalized models from long-tailed data distributions. Many works [zhang2023deep, leevy2018surveyimbalance, johnson2019surveyclassimbalance] have achieved excellent results in class-specific tasks, such as re-sampling, re-margining, and re-weighting [lin2017focal, cui2019classloss, park2021influence, ren2020balanced]. Some studies [qin2023classbalancing, yan2024trainingclassimbalance, um2024self, zhang2024long, katsumata2024labelaug] have extended class-based re-weighting to class-image generation tasks with inter-class distance, balanced distribution, label augmentation, or self-guided methods, showing impressive results. However, since text prompts is a joint distribution of multiple classes and each image cannot be assigned a single class, data balancing in text-image tasks still needs further exploration.
@@ 3 Causal Factors of Concept Composition
In this section, we bridge the gap between synthetic experiments [okawa2023compositionalemerge, wiedemer2023compositionalprinceples, chang2024skews] and text-image generation tasks to further explore the casual factors of concept composition. We conduct controlled experiments on text-image datasets and investigate the influence of crucial factors: model size, dataset scale, and data distribution.
@@
Experiment Setup To ensure dataset the same and training efficiency, we re-caption open-source datasets [laion] and collect a higher quality dataset, as existing open-source pre-trained models [midjourney, sd3, dalle3] do not provide their training data. Our dataset contains 31M text-image pairs with an average text length of 100 words and 20K noun concepts with distributions shown in Figure 2. To ensure the generality of the conclusion, we employ a DiT-based diffusion model [dit] without any special design. To accurately evaluate the model’s concept composition ability rather than clarity, we introduce VQA based on VLM [gpt4v] for success rate as a quantitative metric except the CLIP-Score [clip]. We detect all noun concepts in the caption and pair them, requiring the VLM to verify the existence and relationships of concepts with the following questions: ”Is concept A present in the image?”, ”Is concept B present in the image?” and ”Does the relationship between concepts A and B match the caption?”. When all responses are ”yes,” the sample is considered successful.
Figure 3: The performance of models with different parameter sizes under the LC-Mis benchmark [zhao2024lost].
Model Size In generation tasks, model size also follows the scaling law [kaplan2020scaling], where models with more parameters have stronger generation ability. We keep the VAE from Stable Diffusion [sd3] the same and train diffusion models with 100M, 200M, 500M and 1B parameters on the same 31M dataset from scratch, differing only in the number of blocks and channels. We evaluate the CLIP score and success rate on the LC-Mis benchmark, as shown in Figure 3. When the model size exceeds 200M, the concept composition ability increases at a much slower rate as the model size grows.It indicates that model size is no longer the casual factor for concept composition when it reaches a certain magnitude relative to the dataset.
@@ 4.1 Theoretical Analysis
Without loss of generality, we derive with the $\epsilon$-prediction DDPM [ddpm] framework. It is worth noting that this approach can be easily extended to various other diffusion model variants, such as flow matching [lipman2022flow]. In common implementations [dhariwal2021diffusion, song2020score], the training loss of diffusion models parameterized by $\theta$ can be written as follow:
@@
where $y$ is the text prompts, $x_{t}$ is the noisy image at timestep $t$ from image $x_{0}$ and random Gaussian noises $\epsilon$. Inspired by existing work [somepalli2023diffusionreplication] where image regions respond to prompts as concept phrases, we rewrite the loss function as follow:
@@
On the other hand, we wish that the model responds with similar intensity to different concepts [chefer2023attend]. Thus, we have the response intensity represented by the difference between the conditional and unconditional distribution as follow:
@@
where the intensity depends only on $a_{t}$. Meanwhile, we have the formulation of conditional and unconditional distribution in classifier-free guidance [ho2022classifierfree] as follows:
@@
[…]
@@ 4.2 IMBA Distance
Measuring data balance is particularly challenging due to the complexity of text prompts compared to classes [lin2017focal, qin2023classbalancing]. Additionally, with the exponential growth of datasets, offline data pruning will lead to data waste and significant computational and time costs. In this section, we demonstrate that the IMBA distance is able to represent the frequency proportion of concepts with both synthetic experiments and text-image generation experiments. We further give out the formulation of IMBA distance in Appendix A.
@@ 4.3 IMBA Loss
Compared to offline concept frequency-based loss weights [zhang2023deep, cui2019classloss, park2021influence], IMBA loss is more accurate and efficient. Firstly, since text prompts are the joint distribution of multiple concepts, and each concept has a different frequency, it is challenging to intuitively and quickly set loss weights for samples as in class-related tasks. Secondly, while it is feasible to construct a concept graph where concepts are represented as nodes and the co-occurrence frequency of concept pairs as edge values to derive loss weights, the data distribution that the model learns evolves throughout the training process. This evolution creates a misalignment between the offline loss weights and the model’s understanding. In contrast, IMBA distance is naturally coupled with the training process, providing a more consistent representation of data distribution. Thirdly, since different regions of an image may contain different concepts, data balancing should be performed at the concept region level. IMBA loss supports different loss weights for different regions, making it more accurate compared to a single weight for all regions. Additionally, with the exponential growth of data, constructing an offline concept graph becomes increasingly time-consuming and computationally intensive, making it challenging to reuse across different datasets. IMBA loss, on the other hand, only requires the calculation of IMBA distance with only a few lines of code during training, making it highly efficient and easily reusable. It is worth noting that our method is orthogonal to training-free methods and can be combined to further enhance the concept composition ability.
@@ 4.4 Inert-CompBench
Due to the lack of attention to these concepts in existing benchmarks [huang2023t2icompbench, zhao2024lost], we extract inert concepts from open-world datasets to construct a new benchmark called Inert-CompBench as a supplement. As shown in Algorithm 2, our framework comprises five phases: (1) Extract candidate sets where head concepts exhibit occurrence frequencies exceeding tail concepts by 100:1 ratio based on large-scale dataset statistics. (2) Select n domain-representative entity concepts from each pool through semantic typicality analysis. (3) Construct n×n Cartesian product combination space. (4) Build concept co-occurrence graphs to filter Top-K pairs with minimal structural associations, ensuring test cases reflect non-trivial compositional relationships. (5) Generate 5 linguistically diverse prompts per selected pair using GPT-4, ultimately forming a benchmark with 1K fine-grained test instances. This design forces models to handle statistically weak concept combinations, effectively revealing their compositional reasoning limitations.
@@
3:Select $n$ representative entities: $\hat{\mathcal{H}}\subset\mathcal{H},\hat{\mathcal{T}}\subset\mathcal{T}$ via semantic typicality analysis (Phase 2)
@@
7:Benchmark set $\mathcal{B}=\{\mathcal{P},\mathcal{C}^{*}\}$
@@ 5 Experiments
Setup. To keep the dataset the same and improve training efficiency, we recaption and filter open-source data [laion], resulting in a higher quality dataset of 31M samples for training. We employ a 1B parameter DiT-based diffusion model [dit] without any special design as the pipeline. Text prompts are injected into the diffusion model using T5 model [raffel2020t5], and $512\times 512$ images are encoded into the latent space using the VAE from Stable Diffusion [sd3].
Baseline. We train the diffusion model from scratch for 4 epochs on the new dataset using the original diffusion loss to establish a baseline. To demonstrate the effectiveness of our method, we train the same model for the same epochs on the same dataset with the proposed IMBA loss. Additionally, we implement the training-free method [chefer2023attend] on the baseline for comparison. Furthermore, we finetune the baseline model from 3 epochs for 1 additional epochs with the IMBA loss, demonstrating that our method also provides significant benefits during fine-tuning.
Benchmark. We evaluate the concept composition ability of models on T2I-CompBench [huang2023t2icompbench], LC-Mis [zhao2024lost] and Inert-CompBench. We employ VQA based on VLM [gpt4v] for evaluation in LC-Mis and Inert-CompBench as described in experiments setup of Section 3.
@@
Model LC-Mis [zhao2024lost] T2I-CompBench [huang2023t2icompbench] Inert-CompBench